# Noise-induced resonant acceleration of a charge in an intermittent magnetic field: an exact solution for ergodic and non-ergodic fluctuations

### v3

@@ II Trajectory approach: general exact solution
$\langle r^{2}(t)\rangle=\sum_{n=0}^{\infty}\int_{t_{n-1}}^{t}dt_{n}\int_{t_{n-2}}^{t}dt_{n-1}\dots\int_{0}^{t}dt_{1}\Psi(t-t_{n})\;\;\;\;\;\;$
@@
which thanks to the convolution structure (which is preserved due the exponential time dependence of both $\bf\Sigma$} and $\tilde{M}$) can be exactly summed in Laplace domain. If we now reintroduce the dependence on the initial value of the magnetic field infact this leads to
$\langle r^{2}(s)\rangle^{\pm}=\frac{{\bf\tilde{\Sigma}_{\pm}}(s)\cdot{\it I}+{\bf\tilde{\Sigma}_{\mp}}(s)\cdot\tilde{M}_{\psi_{\pm}}(s)}{1-\tilde{M}_{\psi_{\mp}}(s)\tilde{M}_{\psi_{\pm}}(s)}{\bf\tilde{A}^{\pm}_{0}}$ (25)
where $\tilde{M}_{\psi_{\pm}}(s)$ is the Laplace transform $\mathcal{L}_{s}[\psi(t)\tilde{M}_{\pm}(t)]$ and ${\bf\tilde{\Sigma}_{\pm}}(s)=\mathcal{L}_{s}\left[\int_{0}^{t}d\tau\Psi(t-\tau)\psi(\tau){{\bf\Sigma}_{\pm}}(t-\tau,\tau)\right]$. Eqs. (25) can therefore be expressed as a function only of the Laplace Transform of $\psi(\tau)$, and then used to evaluate $\langle r^{2}(t)\rangle$ via Eq. (16). For Poissonian fluctuations of rate $\gamma$ this leads to
@@ II.2 The symmetric case: $B_{0}=\omega_{0}=0$
Figure 3: $\langle r^{2}(t)\rangle$ for the diffusion of a charged particle in a fluctuating magnetic field $B(t)$, with parameters $\omega_{0}=1$ and $\omega_{1}=0.5$ and for field fluctuations with power-law distribution $\psi(t)_{t\gg 1}\sim(t/\tau)^{-\alpha-1}$ at large times. (Left panel) non-ergodic case of a Mittag-Leffler distribution with parameters $\tau=10$ and $\alpha=1/2$ (diverging mean time). (Right panel) Power-law distribution with finite mean time and diverging second moment, as in Eq. (35) with $\alpha=3/2$ and $\tau=1$. In both panels, the black lines indicate the result obtained from inverting the analytical solution (25). All other lines: numerical simulations. Higher number of runs show closer match with the analytical solution, with slower convergence for the non-ergodic case.
@@ II.3 Slow fluctuations and ergodicity breaking
Fig 3 shows that for these slow fluctuations convergence to the predicted value is slow with increasing number of trajectories, slower in the non-ergodic case. This is due to the fact that many runs tend to be trapped in the same value of the field for long times, not sampling enough phase space.

### v4

@@ II Trajectory approach: general exact solution
$\langle r^{2}(t)\rangle=\sum_{n=1}^{\infty}\int_{t_{n-1}}^{t}dt_{n}\int_{t_{n-2}}^{t}dt_{n-1}\dots\int_{0}^{t}dt_{1}\Psi(t-t_{n})\;\;\;\;\;\;$
@@
to which the contribution $\langle r^{2}(t)\rangle^{\pm}_{0}=\Psi(t){\bf A^{\pm\dagger}_{0}}\cdot\sigma_{\pm}(t){\bf A^{\pm}_{0}}$ from trajectories with no switches of the field has to be added. Thanks to its convolution structure (which is preserved due the exponential time dependence of both $\bf\Sigma$} and $\tilde{M}$) Eq. (’II) can be exactly summed in Laplace domain. If we now reintroduce the dependence on the initial value of the magnetic field infact this leads to
$\langle r^{2}(s)\rangle^{\pm}=\frac{{\bf\tilde{\Sigma}_{\pm}}(s)\cdot{\it I}+{\bf\tilde{\Sigma}_{\mp}}(s)\cdot\tilde{M}_{\psi_{\pm}}(s)}{1-\tilde{M}_{\psi_{\mp}}(s)\tilde{M}_{\psi_{\pm}}(s)}{\bf\tilde{A}^{\pm}_{0}}+\langle r^{2}(s)\rangle^{\pm}_{0}$ (25)
where $\tilde{M}_{\psi_{\pm}}(s)$ is the Laplace transform $\mathcal{L}_{s}[\psi(t)\tilde{M}_{\pm}(t)]$, ${\bf\tilde{\Sigma}_{\pm}}(s)$$\mathcal{L}_{s}\left[\int_{0}^{t}d\tau\Psi(t-\tau)\psi(\tau){{\bf\Sigma}_{\pm}}(t-\tau,\tau)\right]$ and  $\langle r^{2}(s)\rangle^{\pm}_{0}$$\mathcal{L}_{s}\left[\langle r^{2}(t)\rangle^{\pm}_{0}\right].\;$The two terms in the fraction on the right-hand side of Eq. (25) are the result of summing up on trajectories that end either with the field pointing in the opposite direction as at the start (first term), and therefore undergo an even number of field switches before the last one, or pointing in the same direction (second term), and therefore with an odd number of switches before the last one. Eq. (25) can therefore be expressed as a function only of $\mathcal{L}_{s}[\psi(\tau)]$ (or of its translates), and then used to evaluate $\langle r^{2}(t)\rangle$ via Eq. (16). For Poissonian fluctuations of rate $\gamma$ this leads to
@@ II.2 The symmetric case: $B_{0}=\omega_{0}=0$
Figure 3: $\langle r^{2}(t)\rangle$ for the diffusion of a charged particle in a fluctuating magnetic field $B(t)$, with parameters $\omega_{0}=1$ and $\omega_{1}=0.5$ and for field fluctuations with power-law distribution $\psi(t)_{t\gg 1}\sim(t/\tau)^{-\alpha-1}$. (Left panel) non-ergodic case of a Mittag-Leffler distribution with parameters $\tau=10$ and $\alpha=1/2$ (diverging mean time). (Right panel) Power-law distribution with finite mean time and diverging second moment, as in Eq. (35) with $\alpha=3/2$ and $\tau=1$. In both panels, the black lines indicate the result obtained from inverting the analytical solution (25). All other lines: numerical simulations. Higher number of runs show closer match with the analytical solution, with slower convergence for the non-ergodic case.
@@ II.3 Slow fluctuations and ergodicity breaking
Figure 3 shows as well that for these slow fluctuations convergence to the predicted value is slow with increasing number of trajectories, slower in the non-ergodic case. This is due to the fact that many runs get stuck with the same field value for long times, not sampling enough phase space.